OpenStudy (anonymous):

A triangle has a base of 10 centimeters and a height of 7 centimeters. Which is the total area of the triangle? Select one of the options below as your answer:A. 17 square centimeters B. 35 square centimeters C. 70 square centimeters D. 120 square centimeters

OpenStudy (experimentx):

b

OpenStudy (anonymous):

ty

OpenStudy (hoblos):

7*10/2 = 70/2 =35